After extensive medical and marketing research, Pill, Inc., believes it can penetrate the pain reliever market. It is considering two alternative products. The first is to produce a medication for headache pain. The second is a pill for headache and arthritis pain. Both products would be introduced at a price of $9.15 per package in real terms. The headache-only medication is projected to sell 2 million packages a year, while the headache and arthritis remedy would sell 3.5 million packages a year. Cash costs of production in the first year are expected to be $5.05 per package in real terms for the headache-only brand. Production costs are expected to be $5.60 in real terms for the headache and arthritis pill. All prices and costs are expected to rise at the general inflation rate of 2 percent.

Either product requires further investment. The headache-only pill could be produced using equipment costing $18 million. That equipment would last three years and have no resale value. The machinery required to produce the broader remedy would cost $30 million and last three years. The firm expects that equipment to have a $1 million resale value (in real terms) at the end of Year 3.

The company uses straight-line depreciation. The firm faces a corporate tax rate of 34 percent and believes that the appropriate real discount rate is 6 percent.

Calculate the NPV for the Headache-Only product

Calculate the NPV for the Headache and Arthritis Product

-Please show how you got the answer
Answer #1
HEADACHE ONLY MEDICATION: 0 1 2 3
Sales (2000000*9.15+inflation of 2% 18300000 18666000 19039320
Cash production costs (2000000*5.05+inflation of 2%) 10100000 10302000 10508040
Depreciation (18000000/3) 6000000 6000000 6000000
NOI 2200000 2364000 2531280
Tax at 34% 748000 803760 860635
NOPAT 1452000 1560240 1670645
Add: Depreciation 6000000 6000000 6000000
OCF 7452000 7560240 7670645
Capital spending 18000000
FCF -18000000 7452000 7560240 7670645
PVIF at 8.12% [1.06*1.02-1=0.0812] 1 0.92490 0.85544 0.79119
PV at 8.12% -18000000 6892342 6467307 6068953
NPV 1428602
HEADACHE AND ARTHRITIS MEDICATION: 0 1 2 3
Sales (3500000*9.15+inflation of 2% 32025000 32665500 33318810
Cash production costs (3500000*5.60+inflation of 2%) 19600000 19992000 20391840
Depreciation (30000000/3) 10000000 10000000 10000000
NOI 2425000 2673500 2926970
Tax at 34% 824500 908990 995170
NOPAT 1600500 1764510 1931800
Add: Depreciation 10000000 10000000 10000000
OCF 11600500 11764510 11931800
Capital spending 30000000 -700397
FCF -30000000 11600500 11764510 12632197
PVIF at 8.12% [1.06*1.02-1=0.0812] 1 0.92490 0.85544 0.79119
PV at 8.12% -30000000 10729282 10063795 9994494
NPV 787571
NOTE:
TERMINAL NON-OPERATING CASH FLOW OF HEADACHE AND ARTHRITIS PRODUCT:
Sale value in nominal dollar terms = 1000000*1.02^3 = 1061208
Tax at 34% 360811
After tax sale proceeds 700397